### v2

# Top-of-atmosphere radiation over the last millennium reconstructed from proxies

Abstract
Earth’s energy imbalance at the top of the atmosphere is a key climate system metric, but its natural variability is poorly constrained by the short observational record and large uncertainty in coupled climate models. While existing ocean heat content reconstructions offer a longer perspective, they cannot separate the contributions of shortwave and longwave radiation, obscuring the underlying processes. We extend the energy budget record into the pre-industrial period by reconstructing the top-of-atmosphere radiation and related surface variables over the last millennium (850–2000 CE) by using data assimilation to combine proxy data and dynamics from a coupled climate emulator. Validation reveals skill in the reconstructed radiation fields, especially in the tropics. Results show a familiar last-millennium cooling trend, which coincides with persistent heat loss and a reduction in upper-ocean heat content. The cooling trend differs by season and latitude, and is associated with radiative anomalies suggestive of an eastward shift in Indo–Pacific convection. Following large volcanic eruptions, ocean heat content anomalies persist for 10–20 years on average, supporting previous evidence that the cooling trend was forced by decadally-paced eruptions. The reconstruction also reveals that the current rate of energy gain is unprecedented relative to the period before 1850.
1 Introduction
Earth’s energy imbalance (EEI) at the top of the atmosphere (TOA) is a fundamental climate system metric, governing Earth’s total heat content and constraining global temperatures, the hydrological cycle, sea levels, and ice cover (von Schuckmann et al., 2016). Much of the EEI is forced, over the last millennium mostly by volcanic eruptions, solar irradiance, and changes in Earth’s orbit, and more recently by anthropogenic greenhouse gases and aerosols. In addition, there is internal, unforced variability in the energy budget at all timescales. Seasonal variability is primarily due to extratropical storms and the Madden–Julian oscillation, while the El Niño–Southern Oscillation (ENSO) explains most interannual variability (Trenberth et al., 2015). However, decadal to centennial variability in the energy budget remains poorly understood (Wills et al., 2021; Trenberth et al., 2014). This complicates, for example, the interpretation of recent albedo trends (Goessling et al., 2025; Mauritsen et al., 2025; Hodnebrog et al., 2024) and the warming hiatus (Trenberth and Fasullo, 2013; Xie et al., 2016), both of which are not reproduced well by coupled climate models (Olonscheck and Rugenstein, 2024; Raghuraman et al., 2021; Medhaug et al., 2017; Fyfe et al., 2013). A better understanding of natural variability in Earth’s energy budget is also essential to isolating the contribution of anthropogenic forcing in observations, which informs the attribution of historical climate change (Lean, 2018) and helps narrow uncertainty in climate sensitivity (Sherwood et al., 2020).
Our understanding of EEI variability is limited by the short satellite observational record of around 25 years (Loeb et al., 2024), which prohibits the investigation of low-frequency variability and the response to rare, episodic forcings. Even during this well-observed satellite period, discrepancies and uncertainties between EEI estimates are large (Hakuba et al., 2024). Additionally, it is difficult to disentangle forced and internal variability of Earth’s energy budget due to the strong anthropogenic greenhouse gas forcing and the uncertain aerosol forcing during the observed period. Therefore, climate models have been used to investigate aspects of the energy budget like radiative feedbacks (Sherwood et al., 2020) and decadal variability (Palmer and McNeall, 2014; Brown et al., 2014; Zhou et al., 2016), more recently through coordinated experiments like CFMIP and CERESMIP (Webb et al., 2017; Schmidt et al., 2023). However, even coupled models of the CMIP6 generation (Eyring et al., 2016) produce a large spread in TOA radiation, on the order of $5$ to $20\,\mathrm{W\,m^{-2}}$ for most components (Wild, 2020), and systematically underestimate the radiative response to surface warming (Olonscheck and Rugenstein, 2024). In addition, the specified forcings and simulated internal variability may not be accurate (Lean, 2018; Lücke et al., 2023; Fyfe et al., 2021). Other studies have extended the energy budget record by reconstructing the ocean heat content (OHC), which is storing much of the energy gained in recent decades (von Schuckmann et al., 2023). For example, Zanna et al. (2019) and Wu et al. (2025) derive the global-mean OHC over the historical period, while Gebbie and Huybers (2019) reconstruct the gridded OHC over the Common Era. Further OHC reconstructions from the Last Glacial Maximum to present are reviewed in Gebbie (2021). The OHC provides a valuable perspective since it is a more reliable indicator of global change than surface temperatures on interannual to decadal timescales (Palmer and McNeall, 2014; Allison et al., 2020). However, the OHC perspective on the energy budget, particularly in the global mean, obscures the atmospheric processes that cause heat content changes, and the role of clouds and sea ice in mediating them. Rather, a top-of-atmosphere perspective that separates shortwave (SW) and longwave (LW) components is needed to improve the process understanding of energy budget variability.
Here, we reconstruct Earth’s energy budget over the last millennium (850–2000 CE) from paleoclimate proxies. The reconstruction is seasonally and spatially resolved, and partitions the TOA radiation into SW and LW components. We use data assimilation, which combines proxy observations with model dynamics, and has been used successfully to reconstruct past climates (Goosse et al., 2010; Bhend et al., 2012; Steiger et al., 2014, 2018; Perkins and Hakim, 2021; Hakim et al., 2016; Franke et al., 2017; Osman et al., 2021; Erb et al., 2022; Valler et al., 2022, 2024; Judd et al., 2024; Meng et al., 2025; Cooper et al., 2025, e.g.,; see Smerdon et al., 2023 for an overview). Our method exploits the covariance of TOA radiation and OHC with surface temperatures, to which the proxies are sensitive. For example, if a proxy in the Pacific Warm Pool indicates warmer sea surface temperatures (SSTs), the atmospheric response there is to enhance deep convection, which mediates an increase in reflected SW and a decrease in outgoing LW radiation (Fig. 1; e.g., Dong et al., 2019). Similarly, changes in temperature gradients can affect the overturning atmospheric circulation and tropospheric stability, which affects SW radiation through changes in subtropical low clouds (e.g., Mackie et al., 2025; Ceppi and Gregory, 2017; Van Loon and Rugenstein, 2025). Besides clouds, snow/ice cover and solar surface heating play a role, and establish a temperature–radiation relationship over much of the globe (Trenberth et al., 2015). In addition to this physical underpinning, Loeb et al. (2020) demonstrate that the current generation of atmosphere-only climate models can skillfully simulate TOA radiation when SST and sea ice are prescribed; both of these variables can be reconstructed from proxies.
Figure 1: Annual-mean correlation between the SST at a point in the Niño 3.4 region (yellow pentagon; e.g., location of a hypothetical coral proxy) and the TOA (a) reflected SW and (b) outgoing LW radiation globally. Correlations are from last-millennium simulations over 850–1850, averaged over five models using the Fisher z-transformation.
We take three perspectives on the TOA energy balance. The first perspective involves a partition of the EEI into insolation $S$, reflected SW radiation (RSR), and outgoing LW radiation (OLR):
$\text{EEI}^{\downarrow}=S^{\downarrow}-(\text{RSR}^{\uparrow}+\text{OLR}^{\uparrow}).$
By our chosen sign convention, EEI is positive downwelling (energy gain), while RSR and OLR are positive upwelling (energy loss). Separating the SW and LW radiation fields allows us to infer the physical processes governing the EEI, such as clouds.
The second perspective on the TOA energy balance emphasizes the relation to the global-mean surface temperature $T$ in a one-layer energy balance model (Geoffroy et al., 2013; Gregory et al., 2016):
$\text{EEI}=C\frac{dT}{dt}+\gamma T,$ (1)
where EEI is the global-mean TOA energy imbalance, $C$ is the heat capacity of the upper ocean, and $\gamma$ is the heat uptake coefficient of the deep ocean. The product $CT$ is the upper-ocean heat content, which we also reconstruct. On seasonal timescales, such as for the initial response to volcanic cooling, before the upper ocean has adjusted significantly, the deep-ocean heat uptake $\gamma T$ is relatively small. We can then approximately compare OHC and EEI by differentiation or integration. On interannual and longer timescales, the assumption of $\gamma T\ll C(dT/dt)$ breaks down (e.g., Jeevanjee et al., 2025).
A third perspective on EEI involves the forcing–feedback framework (e.g., Sherwood et al., 2015):
$\text{EEI}=R(T)+F_{\text{nat}}+F_{\text{anthro}},$ (2)
where $R(T)$ is the radiative response (also called feedback), and the forcing is separated into natural ($F_{\text{nat}}$; i.e., volcanic and solar) and anthropogenic ($F_{\text{anthro}}$) components. Both forcing and response correlate with temperature-sensitive proxies. Forcing relates to near-surface temperature through its cumulative effect on ocean heat content, and the response through temperature-mediated processes such as Planck and cloud feedbacks (Forster et al., 2021). Thus, the energy imbalance both causes and results from temperature changes. Our method, targeting the pre-industrial last millennium, restricts the reconstruction to the sum of $R(T)+F_{\text{nat}}$, without anthropogenic forcing.
The remainder of the paper is organized as follows. We first describe the data assimilation method (Section 2). After validating this method in pseudoproxy experiments and against instrumental datasets (Section 3), we present our reconstructed energy budget over the last millennium (Section 4), then illustrate applications to the natural energy budget variability and the response to volcanic eruptions (Section 5). Section 6 presents the conclusions.
2 Methods
We use online paleoclimate data assimilation (DA) to reconstruct gridded climate fields at seasonal resolution. DA combines proxy information with model forecasts in a statistically optimal way (e.g., Kalnay et al., 2024). Our DA system comprises three interacting components: (1) linear inverse models (LIMs; Penland and Sardeshmukh, 1995), trained to emulate the last-millennium climate dynamics of CMIP6 models, are used for forecasting; (2) proxy system models (PSMs; Dee et al., 2016; Evans et al., 2013, e.g.,) map from the climate variables predicted by the model to proxy values (e.g., from local SST to coral $\delta^{18}$O); and (3) an ensemble Kalman filter (EnKF; Evensen, 1994) that blends proxy data with model forecasts weighted by their uncertainties, spreading information across space and climate variables. Together, these components provide an estimate of the evolving state of the climate system by continuously updating a model forecast, which maintains memory of the climate from past proxies, with proxy observations during the forecast period.
Specifically, the interaction of the three components proceeds through a forecast–update cycle. A LIM produces a forecast “prior” given an initial state. The first initial state is a random draw from the model climatology, whereas subsequent initial states come from the cycle’s update step described next. The PSMs map this prior state to the expected proxy values, and the EnKF compares these expected values against the actual proxy observations, updating the prior state to produce a posterior estimate. This completes the cycle, which continues in time by using the posterior state as the new initial state to forecast the next season. The methodology for our seasonal-resolution DA is similar to Meng et al. (2025) and based on the prior work by Perkins and Hakim (2021) and Hakim et al. (2016).
2.1 Forecasts using linear inverse models
We use LIMs as efficient emulators of the dynamics and statistics of the last-millennium climate as simulated in five CMIP6 models. Using such emulators allows us to run large ensembles and to incorporate the dynamics of multiple climate models. The LIM dynamics have the form (Penland and Sardeshmukh, 1995; Penland, 1996)
$\frac{d\mathbf{x}}{dt}=\bm{\mathsf{L}}\mathbf{x}+\bm{\mathsf{S}}\bm{\eta},$
where $\mathbf{x}$ is the state vector, $\bm{\mathsf{L}}$ is a matrix representing the linear dynamics, $\bm{\mathsf{S}}$ is the noise amplitude matrix, and $\bm{\eta}$ is a vector of independent Gaussian white noise with unit variance. Together, $\bm{\mathsf{L}}$ and $\bm{\mathsf{S}}$ encode, for example, how surface temperatures and TOA radiation are related. During LIM training, the deterministic dynamics $\bm{\mathsf{L}}$ are derived from a lag-time linear regression, and the noise amplitude $\bm{\mathsf{S}}$ follows from a fluctuation–dissipation relation such that the long-term covariance statistics of the training data are reproduced. During forecasting from one season to the next, each ensemble member starts from a different initial condition and uses a different realization of $\bm{\eta}$. The full equations for determining the LIM matrices and for ensemble forecasts are presented in the Supplemental Information Text S1.
The state vector $\mathbf{x}$ represents seasonal anomalies of atmospheric and oceanic climate variables. During LIM training, the state vector is constructed from last-millennium simulations. During DA, the goal is to estimate the posterior $\mathbf{x}$ by blending the LIM forecast (i.e., the prior $\mathbf{x}$) and proxy information. We include 2-m surface air temperature (SAT), sea surface temperature (SST), TOA energy imbalance (EEI), TOA reflected SW radiation (RSR), TOA outgoing LW radiation (OLR), ocean heat content of the upper 300 m (OHC300), Arctic sea-ice concentration (SIC), and Antarctic SIC. We integrate the OHC over 300 m, which is the depth required to close the seasonal energy budget (Johnson et al., 2023) and is sufficiently deep to remove surface noise (Allison et al., 2020). The OHC also imparts memory to the LIM forecasts since anomalies persist longer than in atmospheric fields.
The LIM training data, consisting of last-millennium simulations over 850–1850, are regridded to the same $\times 2$° grid, converted to anomalies relative to their own 850–1850 climatology, then linearly detrended by season to remove the orbital precession signal. Any low-frequency trend in the reconstruction is therefore driven by the assimilated proxies.
Because the physical gridded state exhibits significant spatial correlation at seasonal resolution, we perform a large dimensionality reduction prior to LIM training. We accomplish this by first subtracting the global means from the gridded state, then calculating the empirical orthogonal functions (EOF) of the area-weighted gridded residuals. Explicitly separating the global means preserves their variance, which is particularly beneficial for spatially noisy fields such as RSR. We then construct the state vector from the global means and truncated principal components (PCs):
$\mathbf{x}=[\tilde{\mathbf{x}}_{1+20}^{\text{SAT}};\tilde{\mathbf{x}}_{1+20}^{\text{SST}};\tilde{\mathbf{x}}_{1+15}^{\text{EEI}};\tilde{\mathbf{x}}_{1+15}^{\text{RSR}};\tilde{\mathbf{x}}_{1+10}^{\text{OLR}};\tilde{\mathbf{x}}_{1+15}^{\text{OHC300}};\tilde{\mathbf{x}}_{10}^{\text{SICn}};\tilde{\mathbf{x}}_{10}^{\text{SICs}}],$
where semicolons denote vertical stacking, and $\tilde{\mathbf{x}}_{1+n}^{\text{var}}$ represents the global mean, followed by the leading $n$ PCs for variable var. For Arctic and Antarctic sea ice (SICn and SICs), we do not separate the global mean and use only the leading 10 PCs. The global mean is standardized by its temporal standard deviation, and the PCs are standardized by the square root of their retained variance after truncation. The truncation ranks $n$ were selected subjectively based on the cumulative variance explained, trading off reconstruction fidelity and degrees of freedom required for LIM training. To ensure a fair comparison across model priors, we use the identical number of PCs for each, despite variations in the actual variance explained (Fig. S1). Mapping this EOF-space state vector $\mathbf{x}$ back to the physical gridded state amounts to a simple linear transformation.
We train LIMs on the CMIP6 last-millennium simulations (“past1000” and “past2k”; Jungclaus et al., 2017) from the following five models: MPI-ESM1-2-LR, CESM2-WACCM-FV2, MRI-ESM2-0, EC-Earth3-Veg-LR, and MIROC-ES2L (Mauritsen et al., 2019; Danabasoglu et al., 2020; Yukimoto et al., 2019; Döscher et al., 2022; Hajima et al., 2020). These are the “model priors,” hereafter referred to as MPI, CESM, MRI, EC-Earth, and MIROC. Last-millennium simulations from other models exist but either lack necessary variables, like SIC, or have not been published. Previous reconstructions like Perkins and Hakim (2021) and Meng et al. (2025) train LIMs on CMIP5 models. However, only since the CMIP6 generation are these models able to faithfully simulate TOA radiation, mainly due to better representations of the low-cloud SW effect (Loeb et al., 2020). We exclude Antarctic sea ice from the MIROC-based LIM during our analysis due to its Antarctic SIC low bias (Hajima et al., 2020).
LIM forecast skill from season to season is important to propagate proxy information through time and to provide accurate priors. Forecast tests show skillful forecast correlations for the global means ($r>0.8$ for all fields at one season lead time; Fig. S2). The LIM tends to be slightly underdispersive, with ensemble spread underestimating forecast error in the ensemble mean (not shown). Spatial skill varies from $r\approx 0.5$ for SAT and OHC300, to $r\approx 0.3$ for radiation fields, with locally better skill in the tropical Pacific. Spatial correlation is primarily lost due to EOF truncation rather than lack of LIM skill, as evident from the reduced correlations of the initial condition. Conversely, the correlations for the global mean are close to unity, highlighting the value of handling it separately in the state vector.
Training on last-millennium simulations means that our reconstruction is most representative for the 850–1850 period. During the historical period (1850 onwards), the energy imbalance is dominated by anthropogenic forcing, for which the relationship with surface temperatures is different than the one learned during the last millennium. Our reconstructed EEI therefore only includes the response and natural forcing, even during the historical period.
2.2 Data assimilation using an ensemble Kalman filter
The Kalman filter is a DA algorithm that optimally combines a prior state, such as a forecast, with available proxy observations, weighted by their uncertainties. In an EnKF, the prior and posterior distributions of the state are approximated as normal distributions using samples (Evensen, 1994). Multiple flavors of the EnKF exist. Here, we use a serial ensemble square root filter (Whitaker and Hamill, 2002).
The EnKF combines the prior ensemble mean $\overline{\mathbf{x}}_{b}$ and a proxy observation $y$ into the posterior ensemble mean $\overline{\mathbf{x}}_{a}$:
$\overline{\mathbf{x}}_{a}=\overline{\mathbf{x}}_{b}+\bm{\mathsf{K}}(y-\bm{\mathsf{H}}\overline{\mathbf{x}}_{b}),$ (3)
where $\bm{\mathsf{K}}$ is the Kalman gain matrix, and $\bm{\mathsf{H}}$ is the forward operator for estimating the observation from the prior. The individual ensemble members are updated such that the posterior covariance (i.e., their spread), estimates the uncertainty consistently with respect to the non-ensemble Kalman filter. The Kalman gain matrix encodes covariances between the observation and the state, which allows us to estimate any field that covaries with surface temperature. The full equations for determining $\bm{\mathsf{K}}$ and for updating the ensemble members are presented in the Supplemental Information Text S2.
Seasonally resolved proxies are assimilated in the season they represent. Assimilating annually resolved proxies is complicated by the fact that the proxy seasonality can span multiple seasons. Our DA strategy in this case is to update all seasons within this time window when the proxy can be estimated by the ensemble (i.e., once the end of the window is reached), similarly to Meng et al. (2025). For example, a proxy with seasonality MAMJJA is assimilated during the JJA step to update the MAM and JJA values. The update to the last season of the window (JJA) then informs future seasons through the LIM forecast, while updates to past seasons are not propagated forward in time. To perform the EnKF update of multiple seasons simultaneously, we use the time-averaging algorithm from Huntley and Hakim (2010). This algorithm uses the time mean over multiple seasons as $\overline{\mathbf{x}}$ in Eq. (3), then adds the deviations of each season around that prior time mean to the posterior time mean.
We perform DA directly on the state vector (i.e., in EOF space) rather than in physical space. This improves computational and storage efficiency because both LIM and EnKF operate in the same low-dimensional EOF space. However, operating in EOF space prevents the use of covariance localization (e.g., Anderson, 2012). Instead, we mitigate sample error by using an ensemble that is much larger than the low-dimensional state vector, and we account for model error by running separate reconstructions with five distinct LIMs. To estimate the proxy value in Eq. (3), the observation operator $\bm{\mathsf{H}}$ must map the state vector to the physical temperature at the proxy location, and subsequently to the proxy value using the PSM (e.g., Hakim et al., 2022).
SIC is physically bounded between zero and one, but the Gaussian distributions of the EnKF do not enforce this constraint. To obtain valid values, we convert SIC anomalies to absolute values by anchoring them to a 1961–1990 climatology, then clip SIC values falling outside the $[0,1]$ interval. The climatology, obtained from Cooper et al. (2025), is based on the multi-model mean from eight CMIP6 historical simulations before 1979, and on satellite observations thereafter.
The LIMs from each of the five model priors are used separately in the DA algorithm, allowing us to sample structural error in the training data, such as differences in covariance relationships. We repeat the reconstruction procedure 20 times for each LIM using a Monte Carlo approach. In each iteration, we assimilate a random sample of 80% of the available proxy data. This allows us to sample over uncertainty in the proxy error estimates by randomly removing proxies that may have an outsized impact (Tardif et al., 2019). The remaining 20% of proxies in each iteration are used for independent validation of reconstruction skill in the pre-instrumental period. Each of the 20 iterations uses 400 ensemble members. We reduce the full 8000-member ensemble per model prior to 400 members by subsampling 20 members from each iteration. The analysis below is performed on the resulting 2000-member multi-model ensemble derived from reconstructions with all five model priors.
2.3 Proxy system models
The EnKF requires PSMs that estimate the proxy values from the current state, corresponding to $\bm{\mathsf{H}}\overline{\mathbf{x}}_{b}$ in Eq. (3). We use linear univariate PSMs that are calibrated over the instrumental period. Such statistical PSMs allow for estimating the proxy error needed for DA from regression residuals, while sacrificing little skill compared to process-based PSMs for most proxy types (Dee et al., 2016; Sanchez et al., 2025). Tree-ring width (TRW) proxies, however, may be poorly modeled by a univariate regression in temperature due to the confounding influence of moisture (Dee et al., 2016). We address this by removing TRW proxies that are primarily moisture-sensitive.
The linear model takes as input the temperature (SAT for terrestrial, SST for marine proxies) at the nearest gridpoint. For seasonally resolved proxies, we fit one PSM for each season. For annually resolved proxies, we objectively determine the proxy seasonality based on the Bayesian information criterion (as in Tardif et al., 2019), then fit a single PSM that takes the mean over those seasons as input. Seasonal DA allows us to explicitly update seasonal temperatures rather than the annual-mean temperature, whereas annual DA can suffer from seasonality biases (Lücke et al., 2021).
The PSMs are calibrated on GISTEMP v4 (Lenssen et al., 2024; GISTEMP Team, 2025) and ERSSTv6 (Huang et al., 2025a, b). We truncate these temperature datasets to the same EOF basis used for each LIM to include representativeness error in the estimated observation error. We remove GISTEMP data before 1900 since there are spatial discontinuities that may affect the calibration. We also remove calibration data after 2000 to avoid the divergence problem, i.e., the decoupling of proxy values from local temperature, such as is common for tree-ring proxies (e.g., D’Arrigo et al., 2008). The full calibration procedure is described in the Supplemental Information Text S3.
2.4 Proxy network
[…]
We average all subseasonal proxies to seasonal resolution, and remove proxies that have a time resolution longer than one year since we cannot assimilate them easily. Proxies are excluded if their overlap with the calibration period is too short (<< 25 years), their calibration correlation is too low (<0.10<0.10), or their error autocorrelation is too high (>0.90>0.90; errors are assumed uncorrelated in time in the Kalman filter). The resulting proxy network comprises 382 records, most of which derive from tree rings (58%) and corals (30%). The proxies have broad spatial coverage, but few are located in the ocean interior, particularly outside the tropics (Fig. S5). The proxy network becomes increasingly sparse toward the beginning of the last millennium, and there are only four seasonally resolved proxies before 1500. Importantly, since annual proxies are sensitive to specific seasons, and the LIM propagates information through proxy-sparse seasons, the algorithm yields seasonal climate information even when there are no seasonal proxies.
2.5 Skill metrics
We validate the reconstruction skill in pseudoproxy experiments, against instrumental datasets, and against withheld proxies. Our primary validation metrics are the Pearson correlation coefficient $r$ and the coefficient of efficiency (Nash and Sutcliffe, 1970)
$\text{CE}=1-\frac{\sum_{i}(v_{i}-x_{i})^{2}}{\sum_{i}(v_{i}-\bar{v})^{2}},$
where $x_{i}$ is the reconstruction time series, $v_{i}$ is the verification time series, and $\bar{v}$ is the time mean of $v_{i}$. Compared to $r$, the CE thus also accounts for differences in signal amplitude and mean, measured relative to the verification variability.
The statistical significance of correlations is determined using the random-phase test, a nonparametric test that generates surrogate time series with the same spectral density (Ebisuzaki, 1997). This accounts for spectral properties such as autocorrelation, which reduces the effective sample size (e.g., Bretherton et al., 1999). For p-values, we use a significance level of $\alpha=0.05$ with the null hypothesis that $r=0$. The alternative hypothesis is $r>0$ if reconstruction skill is compared, and $r\neq 0$ for physical correlations between variables. In some figures and tables, we denote highly significant correlations ($p\leq 0.01$) with **, significant correlations ($p\leq 0.05$) with *, and non-significant correlations ($p>0.05$) with n.s.
3 Validation of reconstruction skill
We validate our method using pseudoproxy experiments (PPEs) and compare the real-proxy reconstruction against instrumental datasets. PPEs help to establish an upper bound on reconstruction fidelity in an idealized scenario (Smerdon, 2012), while instrumental validation assesses the real reconstruction, albeit over a limited time period and against uncertain verification sources (e.g., Thorne et al., 2026; Chan et al., 2025).
3.1 Realistic pseudoproxy experiment
PPEs are particularly useful to assess the feasibility of reconstructing new target variables such as TOA radiation. A climate model simulation serves as “truth,” sampled only by sparse and noisy pseudoproxies that mimic the characteristics of the real proxy network. The challenge is analogous to the real problem, with the benefit that skill can be easily quantified (Smerdon, 2012; Steiger et al., 2014).
Figure 2: Comparison of the pseudoproxy imperfect-model reconstruction with the truth simulation. (a–d) Global-mean time series and their correlation coefficients r over 850–1850. The reconstruction (black) and the truth (yellow) are anomalies relative to 850–1850 with a 20-yr low-pass filter. Radiation after 1850 is dashed since it includes anthropogenic forcing, which we cannot reconstruct. DA is the mean of the four single-model reconstructions with the MPI, CESM, MRI, and EC-Earth model priors. Carets indicate volcanic eruptions, scaled by their volcanic stratospheric sulfur injection, from the eVolv2k (Toohey and Sigl, 2017) and the CMIP7 historical volcanic forcing datasets (Aubry et al., 2025; small eruptions with a VSSI below $0.5\,\mathrm{Tg\ S}$ removed). All correlations of global means are significant as denoted by (**). Shading denotes the 5th–95th percentile range. (e–h) Spatial correlation of annual anomalies over 850–1850 between reconstruction and truth. Correlations that are not significantly positive ($\alpha=0.05$) are hatched. Numbers in the top right corners are global means of the spatial correlation.
For our realistic PPEs, truth consists of the concatenated last-millennium and historical simulations from the MIROC model. We then reconstruct MIROC fields using LIMs trained on the MPI, CESM, MRI, and EC-Earth model priors. This constitutes an imperfect-model experiment, which includes error from a mismatch in LIM dynamics and the EOF basis relative to the truth simulation, analogous to the real reconstruction where the true dynamics are unknown. For each real proxy, a pseudoproxy is drawn from the truth simulation using the real PSM with added noise consistent with the estimated error. The pseudoproxies mimic the real proxies in location, temporal availability, seasonal sensitivity, signal-to-noise ratio, and temperature field sensitivity (SAT or SST).
Results show that the global-mean reconstructions for the fields of interest correlate significantly with the truth simulation (Fig. 2a–d). For SAT, decadal variability is tracked well (annual $r=0.8$), and the truth is generally within the ensemble spread. Modern warming is somewhat underestimated, possibly because patterns of warming are different in the MIROC model compared to the LIMs (Fig. 2a). The reconstruction is also mostly skillful for the TOA radiation fields, although correlations and CE are lower for EEI than for RSR and OLR. Large excursions in TOA radiation coincide with major volcanic eruptions, although the magnitude is underestimated, particularly in the EEI. The variance of the EEI in the PPE is similar to the real reconstruction, suggesting that the PPE may be representative of the real problem. Arctic SIC validates well on sea-ice area and spatial patterns, in contrast to Antarctic SIC, which shows low correlations and a large spread across model priors (Fig. S3). Reconstruction skill for all fields is significantly reduced in the PPE before 1100, when fewer than 70 proxies are available, none of them in the tropical Pacific.
The PPE also demonstrates spatial reconstruction skill over the last millennium (Fig. 2e–h). Correlations with the truth simulation are positive in most regions, although they are smaller over the Southern Ocean, which is relatively sparse in proxies. Like for the global mean, annual skill is highest for SAT and lowest for EEI.
3.2 Pseudoproxy experiment for forcing and response
We run further PPEs to test if our DA system is able to reconstruct both the response and the natural forcing, despite their distinct lead/lag relationships with temperature (e.g., Proistosescu et al., 2018). In these PPEs, we separate the forcing and response in the state vector. The truth and training data come from CMIP6 “hist-nat” simulations (Gillett et al., 2016), with forcing derived from “piClim-histnat” simulations (Pincus et al., 2016), of the CanESM5 (Swart et al., 2019) and NorESM2-LM (Seland et al., 2020) models.
We find skillful reconstructions of the response, natural forcing, and 300-m OHC (Fig. S4, Table S1). Correlations are particularly high for the response and OHC. While forcing skill is relatively high in perfect-model experiments, it is not always significant in imperfect-model experiments. The magnitude of volcanic forcing is greatly underestimated, by up to a factor of five, similar to the realistic PPE experiment (Fig. 2b). The lead/lag relationships of the response and forcing in the models are reproduced by the reconstruction (Fig. S4d), which is critical for proxy information to inform the energy imbalance through DA. Natural forcing also includes solar variability, although its magnitude is below the noise level of the reconstruction. We conclude that reconstructing both response and natural forcing is possible, but with a much reduced magnitude for volcanic forcing. We hypothesize that OHC plays a critical role in linking the energy budget and radiation to temperatures through proxies.
3.3 Instrumental validation of real-proxy reconstruction
Validation against independent datasets allows us to assess the real reconstruction. The SAT, SST, and OHC reconstructions are highly correlated with instrumental products and other reconstructions (Figs. S7 and S9a,b). The Arctic SIC reconstruction is skillful at annual and seasonal scales when compared to satellite observations (Figs. S8 and S9c–f), but Antarctic SIC is not. Further, unassimilated proxies (20% withheld for validation), which are predicted from the reconstruction and compared directly to proxy values, correlate significantly in the network mean, indicating skill over the whole last millennium (Fig. S6).
The Earth Radiation Budget Experiment (ERBE) satellite observations of TOA radiation overlap with our reconstruction over 1985–1999. We validate against the DEEP-C v5 dataset (Liu et al., 2020; Liu and Allan, 2022), which combines these satellite observations with atmospheric reanalyses and atmosphere-only model simulations. We also validate against the ERBE WFOV Edition 4.1 datasets (ERBE Science Team, 2020), which are direct, recently recalibrated observations but lack complete spatial and temporal coverage.
Reconstructed global-mean TOA radiation qualitatively tracks interannual variability in the DEEP-C and ERBE data, with positive but not always significant correlations of $r>0.4$ (Fig. 3a–c). Most of this variability represents the radiative response to internal variability, but there is a forced signal from the 1991 Mt. Pinatubo eruption. As in the PPEs (Section 33.1,3.2), this volcanic forcing is underestimated. Much of the skill comes from the tropics and subtropics, likely mediated by ENSO, but large regions in the mid-latitudes and polar regions also have positive correlations (Fig. 3d–f). The reduced skill at higher latitudes is likely the result of a weaker temperature–radiation relationship, mainly due to weather noise from extratropical storms in TOA radiation (e.g., Trenberth et al., 2015). The reconstructed seasonal TOA radiation is also skillful, but with correlations slightly less than the annual values (Fig. S10). Compared to the PPE, the correlations of the global mean are similar, but the skill mostly comes from the tropics, while it is more uniform in the PPE. The correlations are not robust due to the short overlap period and high temporal autocorrelation.
Figure 3: Annual TOA radiation anomalies from our reconstruction, the DEEP-C combined product, and ERBE satellite measurements. The linear trend over 1985–1999 has been removed. Shading denotes the 5th–95th percentile range. (a–c) Global-mean time series and their correlation coefficients r. (d–f) Spatial correlation of reconstruction with DEEP-C. Numbers in the top right corners are global-mean values. Significance of spatial correlations is not indicated since most are non-significant due to the small effective sample size ($n_{\text{eff}}\approx 10$ to $15$ for global means).
Our reconstruction also correlates well ($r=0.72$, $p\leq 0.01$) with decadal variations in instrumental estimates of the upper-ocean heat content (Fig. 4). For comparison, we use the Minière et al. (2026) compilation of instrumental datasets, derived primarily using in-situ temperature profiles. Most of the decadal variability in OHC appears forced by volcanic eruptions. The recovery of OHC following the 1963 Agung eruption took multiple decades, possibly prolonged by the 1968 Fernandina eruption, unforced decadal variability, and tropospheric aerosol forcing. We conclude that our reconstruction of the energy budget has many skillful aspects, despite the spatiotemporal sparseness of the proxy network and the complex relationship between surface temperature and TOA radiation.
Figure 4: Annual upper-ocean heat content from our reconstruction and the Minière et al. (2026) compilation of instrumental datasets. The linear trend over 1960–2000 has been removed to emphasize interannual to decadal variations as in Church et al. (2005). Only instrumental datasets that span the full period are used, leaving a compilation of eight. Stratospheric aerosol optical depth (SAOD) at 550 nm is from the CMIP7 historical volcanic forcing dataset (Aubry et al., 2025). Shading denotes the 5th–95th percentile range.
4 Reconstruction over the last millennium
Our reconstruction consists of gridded, seasonal climate fields over 850–2000 CE. Results labeled as DA are the combined ensemble of the five single-model reconstructions. Annual means are taken from December to the following November. Uncertainties refer to the very likely range (5th to 95th ensemble percentile, i.e., the 90% median credible interval).
4.1 Global means of temperature and radiation
The reconstructed global-mean SAT has a cooling trend over much of the last millennium, followed by warming over the 1900s (Fig. 5a). The SAT transition from a warm Medieval Climate Anomaly (MCA; c. 800–1200) into a colder Little Ice Age (LIA; c. 1300–1850) is a feature often found in last-millennium proxy reconstructions (e.g., Mann et al., 1999; Esper et al., 2002; Mann et al., 2009; PAGES 2k Consortium, 2019) and simulations (e.g., Fernández-Donado et al., 2013; Otto-Bliesner et al., 2016; Ljungqvist et al., 2019). Based on linear regression over 850–1850, we find a cooling trend of $0.18\,\mathrm{K\,kyr^{-1}}$ (see also Fig. 8a). There is also considerable multidecadal variability, some of which coincides with large volcanic eruptions; we will consider these events in greater detail in Section 44.5. The global-mean SST and OHC300 are highly correlated with SAT ($r>0.8$; OHC shown in Fig. 5d, SST in Fig. S11d). Our reconstruction has a similar mean value, multidecadal variability ($r=0.85$, $p\leq 0.01$ at 20-yr timescales), and trend as the multi-method ensemble from PAGES 2k Consortium (2019). This is also the case when comparing to LMR v2.1 (Tardif et al., 2019) and PHYDA (Steiger et al., 2018) in Fig. S11a. The NH extratropical land summer temperatures are at the cold end of comparable reconstructions (e.g., Wilson et al., 2016; Büntgen et al., 2021), and colder than instrumental temperatures before 1950, while the SH temperature agrees well with Neukom et al. (2014), as shown in Fig. S11b,c.
Figure 5: Global-mean time series over the last millennium. DA refers to the combined ensemble of the multi-model reconstructions. Shading denotes the 5th–95th percentile range. Bold lines have a 20-yr low-pass filter; thin, faint lines are annual values. (a) SAT anomalies relative to 1961–1990, comparing to the multi-method ensemble from PAGES 2k Consortium (2019) and the HadCRUT5 instrumental dataset (Morice et al., 2021). (b) Absolute energy imbalance (rather than anomalies), representing the radiative response and natural forcing. The orange line shows the rate of OHC change from Gebbie and Huybers (2019), expressed as equivalent EEI and divided by 90%, which is the fraction of energy imbalance absorbed by the ocean (von Schuckmann et al., 2023). Gebbie and Huybers (2019) data after 1850 is dashed since it includes anthropogenic forcing, which we cannot reconstruct. Carets indicate volcanic eruptions, scaled by their volcanic stratospheric sulfur injection, from the eVolv2k (Toohey and Sigl, 2017) and the CMIP7 historical volcanic forcing datasets (Aubry et al., 2025; small eruptions with a VSSI below $0.5\,\mathrm{Tg\ S}$ removed). Those marked in orange are composited in Fig. 10. (c) RSR and OLR, constituting the EEI in (b). (d) Upper-ocean heat content anomalies, compared to the reconstructions by Gebbie and Huybers (2019) and Wu et al. (2025).
Figure 6: Recent energy imbalance in the context of pre-industrial variability. The EEI reconstructed in this study is compared to CERES EBAF Ed4.2.1 data (Doelling, 2025; Loeb et al., 2018), and to the 20-yr sliding window trend in the heat content reconstructions of IGCC (Forster et al., 2025; Smith et al., 2025) and Wu et al. (2025). The OHC from Wu et al. (2025) has been divided by 90%, which is the fraction of energy imbalance absorbed by the ocean (von Schuckmann et al., 2023). Shading denotes the 5th–95th percentile range, assuming independent errors when adding or subtracting time series. (a) Energy imbalance trends over 25-yr periods, which is the number of full years in the CERES record. The histogram shows the distribution of 25-yr trends in the ensemble-mean EEI from our reconstruction over 850–1850. The yellow line shows the kernel density estimate of 25-yr trends in the five CMIP6 last-millennium simulations. The trend from CERES EBAF is shown in blue, with an uncertainty of $\pm 0.20\,\mathrm{W\,m^{-2}\,dec^{-1}}$ as in Raghuraman et al. (2021). (b) Absolute energy imbalance over 20-yr periods. The black line is the sum of the reconstructed EEI and the anthropogenic forcing from IGCC. The vertical histogram shows the distribution of 20-yr-mean, ensemble-mean EEI from our reconstruction over 850–1850, and dotted lines show its minimum and maximum. (c) Response and natural forcing over 20-yr periods. The yellow line is the EEI from Wu et al. (2025) minus the anthropogenic forcing from IGCC, to which we anchor our EEI anomalies in order to obtain absolute values.
The absolute EEI governs changes in the energy budget (e.g., a positive EEI implies energy gain), whereas the reconstruction provides anomalies. To obtain absolute values, we anchor the reconstructed EEI anomalies, which contain the response and natural forcing, to an independent climatology of the total EEI minus the anthropogenic forcing, based on 20-yr running means with centers during 1880–1960. The equivalence follows from rearranging Eq. (2):
$\underbrace{R+F_{\text{nat}}}_{\text{DA}}=\underbrace{\text{EEI}}_{\text{Wu et al. (2025)}}-\underbrace{F_{\text{anthro}}}_{\text{IGCC}},$
where the EEI from Wu et al. (2025) is the rate of change of reconstructed ocean heat content, and the anthropogenic forcing from IGCC (Indicators of Global Climate Change, an update to metrics from the IPCC AR6 report; Forster et al., 2025; Smith et al., 2025) is estimated using empirical relationships. Our anomalies and the anchor timeseries have similar variability over the anchoring period of 1880–1960 (Fig. 6c).
The cooling trend over the last millennium coincides with energy loss (Fig. 5b), manifesting as a negative energy imbalance over 850–1500. After that, the energy imbalance is close to zero. The only other proxy-based estimate of the last-millennium energy budget is the OHC reconstruction from Gebbie and Huybers (2019, version OPT-0015). At multidecadal timescales, our reconstruction has a weakly negative correlation (r=0.10r=-0.10, $p=0.73$ at 50-yr timescales over 850–1850) with the Gebbie and Huybers (2019) estimate (Fig. S12a). The disagreement increases before 1100, although our reconstruction is likely less reliable during this early period, as discussed in Section 33.1. Over 1100–1850, our mean imbalance ($0.13\pm 0.13\,\mathrm{W\,m^{-2}}$) agrees well with theirs ($0.12\,\mathrm{W\,m^{-2}}$), and consequently, the integrated energy imbalance also agrees well (Fig. S12b), indicating a loss of about $1700\,\mathrm{ZJ}$ over 850–1850. After 1850, our reconstructions diverge because Gebbie and Huybers (2019)’s estimate also includes anthropogenic forcing. Further, their prescribed SSTs show stronger and earlier warming than ours (Figs. 5d and S11d).
The global-mean RSR and OLR are dominated by multidecadal variability, often associated with volcanic eruptions (Fig. 5c). RSR is anticorrelated with SAT (annual r=0.78r=-0.78, $p\leq 0.01$), while OLR has a strong positive correlation (annual $r=0.97$, $p\leq 0.01$). These correlations between temperature and TOA radiation are about $20\%$ stronger than in last-millennium simulations, likely since our method depends on temperature to reconstruct radiation. In contrast, the correlation of EEI with SAT is weakly negative (annual r=0.21r=-0.21, $p=0.02$), while it is weakly positive in models. RSR has more variance than OLR at all timescales.
4.2 Context for recent energy imbalance
We consider the recent energy imbalance and its trend in the context of pre-industrial, natural variability. The EEI trend over 25-yr periods during 850–1850 in our reconstruction ranges from 0.26-0.26 to $0.32\,\mathrm{W\,m^{-2}\,dec^{-1}}$ (Fig. 6a). The distribution of these pre-industrial EEI trends provides context for the modern EEI trend. The trend over the 25-yr CERES satellite period (2001–2025) as provided in CERES EBAF data (Doelling, 2025; Loeb et al., 2018) is $0.44\,\mathrm{W\,m^{-2}\,dec^{-1}}$, far exceeding pre-industrial values. Even a trend of $0.24\,\mathrm{W\,m^{-2}\,dec^{-1}}$, corresponding to the lower uncertainty bound of the CERES EBAF trend, is in the 99th percentile of pre-industrial variability in our reconstruction. Last-millennium simulations have a slightly wider range of EEI variability, likely due to a stronger response to volcanic eruption. The CERES trend of $0.44\,\mathrm{W\,m^{-2}\,dec^{-1}}$ is in the 97th percentile of pre-industrial variability from last-millennium simulations.
EEI ranges from 0.41-0.41 to $0.07\,\mathrm{W\,m^{-2}}$ over 20-yr periods during 850–1850 in our ensemble-mean reconstruction (Fig. 6b). We compare this pre-industrial distribution to the modern EEI estimated from CERES EBAF satellite data, the IGCC heat inventory (Forster et al., 2025; Smith et al., 2025), and the OHC reconstruction by Wu et al. (2025). In these three modern datasets, the EEI over all 20-yr periods starting after 1915 exceeds any pre-industrial 20-yr-average EEI, although the EEI around 1960 almost falls within the pre-industrial distribution. Considering not just the distribution of the ensemble mean but also all ensemble members, the maximum pre-industrial EEI is $0.82\,\mathrm{W\,m^{-2}}$. With this more conservative threshold, the EEI for all 20-yr periods starting after 1990 is unprecedented in the pre-industrial period.
While our reconstruction does not contain the anthropogenic forcing, we can add its estimate from IGCC to our reconstructed EEI. The sum then agrees well with the total EEI from Wu et al. (2025) over 1880–1960, although there appears to be a 10-yr lag (Fig. 6b). After 1960, the response in our reconstruction is likely underestimated (Fig. 6c).
4.3 Sea ice
Sea ice influences high-latitude TOA radiation through surface albedo, lapse rate, and cloud feedbacks (Jenkins and Dai, 2021). We therefore evaluate Arctic and Antarctic sea-ice area (SIA), calculated as the sum of the products of SIC and grid-cell area in each hemisphere (e.g., Notz, 2014). While Arctic sea-ice reconstructions over the last millennium exist (e.g., Brennan and Hakim, 2022; Meng et al., 2025), Antarctic sea ice has so far been elusive (Thomas et al., 2019), and we caution that our Antarctic reconstruction shows much lower skill than the Arctic based on pseudoproxy experiments (Fig. S3) and instrumental validation (Figs. S8 and S9).
Figure 7: Annual-mean sea-ice area over the last millennium. Shading denotes the 5th–95th percentile range. We show anomalies relative to 1961–1990 with 20-yr low-pass filter and compare to the reconstructions from Brennan and Hakim (2022) for the Arctic and Dalaiden et al. (2023) for the Antarctic.
SIA in both hemispheres increased with the cooling trend over the last millennium (Fig. 7). Arctic SIA agrees closely with Brennan and Hakim (2022), which is a proxy-based reconstruction using offline DA (no model forecasts). Over 850–1850, the annual-mean Arctic SIA grows by $0.3\times 10^{6}\,\mathrm{km^{2}}$, mostly in the Barents Sea, while $0.9\times 10^{6}\,\mathrm{km^{2}}$ are lost over 1850–2000 (Figs. S8 and S13). The peak SIA during the LIA is slightly lower than in Brennan and Hakim (2022). Antarctic SIA has less variability and more uncertainty across model priors when compared to Arctic SIA. Interannual to decadal variability in Antarctic SIA over 1700–2000 correlates significantly with Dalaiden et al. (2023), another proxy-based reconstruction using offline DA. There is no clear trend in Antarctic sea-ice loss over 1850–2000 (Fig. S8) in our reconstruction, Cooper et al. (2025), and Dalaiden et al. (2023).
4.4 Last-millennium trends
The global-mean cooling trend from the MCA into the LIA is a prominent feature in our reconstruction. Associated with it are regional variations, which we summarize here in terms of zonal-mean trends, based on a linear regression over 850–1850; spatial trends are shown in Fig. S14.
Figure 8: Zonal-mean trends over 850–1850. Each panel shows (top) the annual mean and (bottom) the seasonal means. Yellow lines correspond to the CMIP6 last-millennium simulations from the MPI, CESM, MRI, EC-Earth, and MIROC models. Numbers denote the global-mean trend (reconstruction in black with 5th–95th percentile range, models in yellow with min–max range). Shading denotes the 5th–95th percentile range.
Figure 9: Seasonal high-latitude (>> 60°) temperature trends over the last millennium for the Arctic and Antarctic. Trends are expressed as departures in seasonal temperature trends from the annual-mean trend over 850–1850. The high-latitude (>> 60°) insolation trend (yellow) from Berger and Loutre (1991) is shown as daily (dotted) and seasonal-mean (solid) values. Shading denotes the 5th–95th percentile range.
The global-mean cooling trend over 850–1850 is $0.18\,\mathrm{K\,kyr^{-1}}$, with a 5th–95th percentile range of 0.29-0.29 to $0.06\,\mathrm{K\,kyr^{-1}}$ (Fig. 8a). In the multi-method reconstruction from PAGES 2k Consortium (2019), the mean cooling trend is $0.30\,\mathrm{K\,kyr^{-1}}$, with a 5th–95th percentile range of 0.51-0.51 to $0.03\,\mathrm{K\,kyr^{-1}}$. In comparison, this trend in the five CMIP6 last-millennium simulations ranges from 0.11-0.11 to $0.28\,\mathrm{K\,kyr^{-1}}$. Coupled models thus tend to simulate last-millennium weak cooling or warming, contradicting proxy evidence. Cooling is stronger in the Northern Hemisphere (NH; $0.22\,\mathrm{K\,kyr^{-1}}$) than in the Southern Hemisphere (SH; $0.14\,\mathrm{K\,kyr^{-1}}$) in our reconstruction. While almost all regions cool from the MCA into the LIA, polar amplification is evident in the SAT trend. Arctic amplification, measured by the ratio of the 67°N–90°N trend to the global-mean trend in SAT, is around $4.4$, about $15\%$ higher than in recent decades (Rantanen et al., 2022).
Trends in SST and OHC are closely related to each other (Fig. 8b,c), with cooling at all latitudes, particularly in the Pacific and North Atlantic. However, there is significant uncertainty in the magnitude of reconstructed cooling in the northern mid-latitudes across model priors. Warming occurs around the Kuroshio–Oyashio Extension and the Gulf Stream (Fig. S14b).
Trends in TOA radiation have a more complex spatial structure that varies by season (Figs. 8d–f). The OLR trend is strongly negative and spatially defined by the Indo–Pacific convective region, whereas the RSR trend is positive and has large local trends at all latitudes. Together, this makes the global-mean EEI trend positive (toward energy gain), until the absolute EEI is approximately zero on centennial timescales after 1500. The RSR and EEI trends are not significantly different from zero, and vary strongly by location and season (Fig. S14d–f).
At high latitudes (>> 60°), seasonal differences in temperature trends are most pronounced (Fig. 9). The Arctic cooling trend is most enhanced in DJF and most subdued in JJA, which increases the amplitude of the seasonal cycle in Arctic temperatures during the last millennium. In contrast, Antarctic cooling is most enhanced in SON and most subdued in MAM and JJA, reducing the seasonal cycle amplitude. The seasonal temperature trends are compared with high-latitude insolation trends, which differ between the hemispheres. Extremes of insolation trends are smoothed in the seasonal averages, which obscures lead–lag relationships at seasonal timescales. However, a lag of up to one season is evident for the Arctic.
4.5 Volcanic eruptions
We composite pre-industrial tropical eruptions with a volcanic stratospheric sulfur injection (VSSI) of at least $9\,\mathrm{Tg\ S}$ (similar to Mt. Pinatubo in 1991) and a latitude equatorward of 30°. For volcanic events within 10 years of each other, we only use the second eruption. Eruption dates, latitude, and VSSI are taken from the eVolv2k version 4 database (Toohey and Sigl, 2017; Sigl and Toohey, 2024), which is derived from ice core records.
Figure 10: Composite analysis (also known as superposed epoch analysis) of global-mean seasonal anomalies for large tropical volcanic eruptions over 850–1850 (n = 14). Eruptions have a VSSI of at least $9\,\mathrm{Tg\ S}$, a latitude equatorward of 30°, and are marked in orange in Fig. 5a. Anomalies are computed relative to the 5 years prior to each eruption, and only the second eruption is used for volcanic double events. The green line in (b) corresponds to the 1695 eruption. The orange line in (d) is the composite of the time derivative of OHC300, rescaled by Earth’s ocean fraction (71%) to obtain the equivalent EEI and smoothed with an annual running mean. This is expected to differ from the time derivative of the composite of OHC300 in (b) since the pre-eruption climatology is different.
There is a clear global cooling signal, albeit with large scatter across individual eruptions, peaking one to two years after the eruption at $0.17\,^{\circ}\mathrm{C}$ (NH: $0.24\,^{\circ}\mathrm{C}$, SH: $0.12\,^{\circ}\mathrm{C}$) on average and fully recovering after 10–15 years (Fig. 10). The stronger NH cooling, particularly in summer (Fig. S15a), is possibly due to proxy distribution and seasonality, but model simulations also suggest a hemispherically asymmetric response (Pauling et al., 2021). Cooling is within the spread of the proxy reconstruction and model composite by He et al. (2026). OHC300 is reduced by around $25\,\mathrm{ZJ}$ on average, most rapidly during the first two years after the eruption. This heat loss is coincident with enhanced RSR (up to $0.8\,\mathrm{W\,m^{-2}}$), a small reduction in OLR (up to $0.6\,\mathrm{W\,m^{-2}}$) and negative EEI of up to $0.2\,\mathrm{W\,m^{-2}}$. EEI becomes positive after 3–5 years, and the OHC recovers over the course of a decade. Sea-ice expansion has a temporal response similar to surface temperature, and there is no discernible impact on ENSO (Fig. S15), consistent with other reconstructions and models (He et al., 2026; Zhu et al., 2022; Dee et al., 2020; Dee and Steiger, 2022). Compared to coupled-model simulations, the volcanic response in all fields is muted and time-lagged (Fig. S16). Relative to the reconstruction, model-simulated cooling is three times larger, and tTOA radiation anomalies are 10–20 times larger.
Multiple large eruptions spaced a few years apart can initiate periods of prolonged OHC loss (Fig. S17; Gupta and Marshall, 2018). For example, the 1693 and 1695 eruptions caused a decade of OHC loss, peaking at $70\,\mathrm{ZJ}$ (green line in Fig. 10b). The recovery from these two eruptions took 30 years due to five additional, smaller eruptions.
4.6 Consistency between OHC and EEI
Approximately 90% of the heat absorbed due to a non-zero EEI is stored in the ocean (von Schuckmann et al., 2023). Recently, 42% of the accumulated OHC has been stored in the upper 300 m (Pan et al., 2025). Therefore, a strong correlation between the EEI and the rate of change of the upper-ocean heat content (dOHC300/dt) is expected, particularly at seasonal to annual timescales, as described by Eq. (1). This can provide insight into the connection between Earth’s energy budget at the TOA and at the surface, but also serves as a consistency check for the DA procedure. We multiply dOHC/dt by Earth’s ocean fraction (71%) to obtain the equivalent EEI over the total surface area. The derivative is calculated using a central time difference as in Trenberth et al. (2014).
Reconstructed EEI and dOHC300/dt are correlated annually at $r=0.64$ ($p\leq 0.01$; Fig. S18). The volcanic response is similar in shape and timescale, but EEI is weaker than dOHC300/dt (Fig. 10d). At seasonal timescales, the correlation is $r=0.50$ ($p\leq 0.01$), but dOHC/dt leads EEI by about one season. With increasing timescale beyond four years, the two quantities decorrelate (Fig. S18e).
5 Discussion
Our reconstruction of Earth’s energy budget over the last millennium represents a new application of paleoclimate DA. After establishing the reconstruction skill and its limitations, this new dataset allows us to consider several aspects of natural energy budget variability, constrained by proxy information and model dynamics.
5.1 Reconstruction skill and limitations
The reconstruction presented here is the first proxy-constrained dataset of TOA radiation over the last millennium. We use the fact that radiation covaries with surface temperature, to which the proxies are sensitive. However, the proxy information is temporally and spatially sparse, and the link to TOA radiation is mediated by complex and noisy physical processes. Therefore, we rely on pseudoproxy experiments, instrumental validation, and internal consistency to establish TOA radiation reconstruction skill. As in previous reconstructions, surface variables generally agree well with instrumental products and other reconstructions (Figs. S7–S9 and S11), and withheld proxies for independent validation over the whole last millennium have a significantly positive network-mean correlation (Fig. S6). However, Antarctic SIC is likely unreliable, and NH extratropical land summer temperatures are too cold.
The energy imbalance at interannual timescales can be reliably reconstructed. This follows from the PPEs (Figs. 2 and S4) and the satellite comparison (Fig. 3). Interannual variability is dominated by the radiative response to unforced surface temperature variability (Trenberth et al., 2015), to which our proxies are sensitive. Indeed, our PPEs showed that the radiative response can be reconstructed accurately from proxies (Fig. S4 and Table S1), explaining the high interannual EEI skill. The magnitude of volcanic forcing, another important but sporadic source of interannual variability, however, is greatly underestimated. This is the case both in PPEs (Figs. 2 and S4) and in the real reconstruction (Fig. 10). Nonetheless, a volcanic signal is present and the relationship to other variables is physically consistent. In particular, our DA system is able to capture the complex lead/lag relationship of forcing and response with temperature (Fig. S4d; Proistosescu et al., 2018).
The constituent RSR and OLR of the energy imbalance have good global-mean skill in the PPE (Fig. 2). In the satellite comparison, correlations are not significant, although this is very likely a result of the short overlap period (Fig. 3). Spatial OLR skill is higher than spatial RSR skill. This is not surprising since OLR covaries strongly with SAT through the Planck response and is spatially coherent, especially in association with ENSO and other modes of internal variability. RSR has slightly lower correlations because it is generally noisier at seasonal timescales due to extratropical storms and the Madden–Julian oscillation (e.g., Trenberth et al., 2015). In the PPE, the EEI skill is lower than RSR and OLR, possibly because the EEI is the small residual of two large quantities.
Consistency between the TOA radiation and OHC, even though not explicitly enforced by the DA system, is further evidence of reconstruction skill (Figs. S18 and 10d). EEI and dOHC300/dt should correspond since much of the energy imbalance is absorbed into the ocean (von Schuckmann et al., 2023; Palmer and McNeall, 2014; Allison et al., 2020). Indeed, their seasonal and annual correlation are moderately high ($r=0.5$ to $0.6$), with a correlation peak at a timescale of 4 years ($r=0.7$). The correlations between EEI and dOHC/dt are similar for instrumental estimates, and would likely be higher if atmospheric heat storage were considered (Hakuba et al., 2024; Mayer et al., 2026). Up to timescales of 20 years, the correlation is above $0.6$, with a trend toward decorrelation for longer timescales. We speculate that the multidecadal correlation might be higher if OHC were to include the deeper ocean. At seasonal timescales, the EEI lags dOHC/dt by one season (Fig. S18d). This may be a result of the more direct covariance of OHC with proxies, while EEI mostly follows from LIM dynamics, or an indication that much of the interannual variability in EEI results from internal fluctuations in OHC rather than external forcing (cf. Proistosescu et al., 2018).
One caveat of our method is that we can only reconstruct the radiative response and natural forcing, but not the anthropogenic forcing dominating the historical period. This is because the temperature–radiation relationship in the LIM is learned from last-millennium simulations, which do not include a strong greenhouse gas forcing signal since they end in 1850. A nonlinear reconstruction framework may be required to capture relationships that change over time and with forcing type. However, if we add an estimate of the anthropogenic forcing to our reconstructed EEI, the sum agrees well in magnitude and phasing with an OHC-based estimate of the total EEI (Fig. 6). Only after 1960 do they diverge as a result of an underestimated radiative response in our reconstruction.
The reconstruction skill diminishes as proxies become more sparse toward the early part of the last millennium. Additionally, there may be a drop in skill before 1100, when no coral proxies are available to constrain the tropical Pacific. This drop is evident in PPEs but not in the withheld-proxy correlations. Further, our reconstruction has a last-millennium warming trend over North America and parts of continental Asia, particularly in summer (Fig. S14a). This trend may be spurious and would explain the negative correlations for some TRW proxies before 1600 (Fig. S6d). During the development of our DA system, we also experimented with other PSM calibration datasets and periods. While the presented version agrees best with instrumental temperatures, other versions tended to have more temperature variance, resulting in a colder LIA and a steeper last-millennium cooling trend. We suggest that PSM calibration uncertainties should be considered in future DA reconstructions, for example, by sampling calibration datasets and periods across Monte Carlo iterations, and by making use of recent instrumental temperature ensembles (e.g., Chan et al., 2024; Lenssen et al., 2024).
We conclude that the reconstruction is skillful for surface fields, the radiative response associated with internal variability, and the timing of volcanic eruptions, despite the spatiotemporal sparseness of the proxy network and the complex relationship between temperature and radiation. However, the magnitude of volcanic forcing is greatly underestimated, and anthropogenic forcing is not represented. Most reliable are global means at interannual to decadal timescales after 1100, although even seasonal fields have considerable skill. The reconstruction fidelity is highest in the tropics and subtropics.
5.2 Earth’s energy budget over the last millennium
Reconstructions and CMIP5 models agree that there was a global-mean cooling trend over the last millennium, although the patterns of cooling may not have been globally synchronous (PAGES 2k Consortium, 2019; Atwood et al., 2016; Neukom et al., 2019). This cooling was driven primarily by volcanic forcing, with a potential minor role of greenhouse gas and solar forcing (Schurer et al., 2014; Büntgen et al., 2020; Wanner et al., 2022; Esper et al., 2012), then amplified by positive feedbacks (Atwood et al., 2016). Since the climate system is overall stable due to the strongly negative Planck feedback, it responds to cooling with reduced OLR. Indeed, the OLR decreased by $0.35\,\mathrm{W\,m^{-2}\,kyr^{-1}}$, while the RSR only increased by $0.16\,\mathrm{W\,m^{-2}\,kyr^{-1}}$, which explains the increasing EEI trend of $0.18\pm 0.21\,\mathrm{W\,m^{-2}\,kyr^{-1}}$, although it is not significantly different from zero (Fig. 8). To sustain the cooling trend, the EEI needs to remain negative on average despite the increasing trend, leading to a slowdown of cooling over centuries. Indeed, both our reconstruction and the OHC reconstruction of Gebbie and Huybers (2019) have a 1100–1850 mean EEI of $0.1\,\mathrm{W\,m^{-2}}$, and a total integrated energy loss over 850–1850 of $1600$ to $1800\,\mathrm{ZJ}$ (Fig. S12). For comparison, Earth has gained about $600\,\mathrm{ZJ}$ over the much shorter 1880–2020 period (Wu et al., 2025).
While the anthropogenic origin of the modern global warming is well-established (Masson-Delmotte et al., 2021), it is merely a “symptom” of a positive energy imbalance (von Schuckmann et al., 2016). By extending the record of energy budget variability to 850, our reconstruction reveals that the current rate of energy gain, which is the fundamental driver of global climate change (Mauritsen et al., 2025; von Schuckmann et al., 2016; Trenberth et al., 2014), is unprecedented in the last millennium. Specifically, the energy imbalance over any 20-yr period after 1990, and possibly already after 1915, exceeds any 20-yr average over 850–1850. Moreover, the 2001–2025 energy imbalance trend is unprecedented among all 25-yr EEI trends in our reconstruction, and is in the 97th percentile of last-millennium model variability (Section 44.2).
By comparing the spatial structure of the radiation trends (Fig. S14d–f), we can speculate on the physical processes involved. The OLR trend is negative almost everywhere, presumably due to the Planck feedback. Superimposed is an El Niño-like response, with an eastward shift of the Indo–Pacific convection region toward the Central Pacific (Hartmann, 2016), causing negative OLR trends in regions with enhanced deep convection and positive OLR trend over the West Pacific Warm Pool. The shift is likely driven by west Pacific cooling, which moves the ascending branch of the Walker circulation to the east (Fig. S14b). In addition to this shift, the EEI has trends in the subtropical east Pacific, likely associated with the subtropical stratus decks. There is also a clear-sky RSR contribution from anomalous Antarctic sea-ice area (Fig. S13a).
Gebbie and Huybers (2019) provide the only other proxy-based reconstruction of the last-millennium energy budget. They propagate SSTs into the ocean interior using a constant circulation model. Annual-mean proxy SSTs from the Ocean2k database (McGregor et al., 2015) are used before 1870, and instrumental winter SSTs after. The SSTs are prescribed as global-mean values rather than spatially resolved fields, but the reconstruction is constrained by basin-wide averages of ship-based vertical temperature profiles in the 1870s. Their SSTs are systematically colder over the last millennium and start to warm about 150 years before most instrumental datasets and our reconstructed SSTs (in 1750 rather than 1900; Fig. S11d), although recent corrections to instrumental data may support a colder 19th century (Chan et al., 2024, 2025). The cooler SSTs also imply a lower OHC (Fig. 5d), but the pre-1850 energy imbalance is not affected by such a mean offset. Our EEI reconstruction is not significantly correlated with theirs, but the mean EEI is similar (Fig. S12). The lack of correlation and the offset in pre-industrial SSTs may be due to differences in proxy network (e.g., the lack of tropical Pacific proxies in Ocean2k), the prescription of global-mean rather than spatially resolved SSTs, the method of proxy-to-temperature conversion, and the blending of winter and annual SSTs. Ocean2k, on which Gebbie and Huybers (2019) is based, may not provide a suitable dataset for annual values in temperature units, rather than the 200-yr-binned standard deviations of proxy units presented in McGregor et al. (2015). However, by resolving the deep ocean, Gebbie and Huybers (2019) may be able to better constrain the energy imbalance on long timescales.
5.3 Seasonality in last-millennium trends
The last-millennium trends of all fields differ consistently by season. The seasonal cycle in the global-mean cooling trend ranges from $0.16\,\mathrm{K\,kyr^{-1}}$ in MAM and JJA to $0.19\,\mathrm{K\,kyr^{-1}}$ in SON and DJF (Fig. 8). High-latitude temperature and insolation trends have a lag correlation, with higher insolation corresponding with warmer temperatures (Fig. 9). The lag of around one season may be a reconstruction artifact, but it could also be a real property of the high-latitude climate, e.g., due to thermal inertia. In the Arctic seasonal cycle, the annual maximum temperature is nearly synchronous with maximum insolation, but the annual minimum temperature lags insolation by around two months (Donohoe et al., 2020).
While the annual-mean insolation is approximately constant over the last millennium, there are latitude-dependent seasonal trends in insolation due to axial precession (cf. Fig. 2d in Lücke et al., 2021). Proxies will reflect these seasonal trends based on their location and seasonality; seasonal DA explicitly accounting for this fact is therefore is essential to separating seasonal and annual-mean trends.
5.4 Climate response to volcanic eruptions
Large volcanic eruptions are the dominant pre-industrial climate forcing (Büntgen et al., 2020). Clusters of them are thought to be the main cause of the LIA (Miller et al., 2012; Brönnimann et al., 2019; Gupta and Marshall, 2018) and previous cold periods (Van Dijk et al., 2024). Sea-surface cooling after eruptions is initially damped by the upper ocean and absorbed into the deeper ocean (Gregory et al., 2016). The cold anomalies at depth re-emerge over subsequent years, extending surface cooling well beyond the volcanic aerosol residence time of 1–2 years. In our volcanic composite, OHC does not recover until 10–20 years after the eruption (Fig. 10b). A decadal pacing of large eruptions would thus allow heat loss to accumulate and persist (Gupta and Marshall, 2018; Zhong et al., 2011). Our results show OHC anomalies of up to $80\,\mathrm{ZJ}$ for volcanic clusters during the late 1200s, late 1400s, early 1700s, and early 1800s (Fig. S17). Miller et al. (2012) hypothesize that the late 1200s cluster (annotated in Fig. 5b), with major eruptions in 1276 and 1286, initiated the cooling trend leading to the LIA. Indeed, the 20-yr period with the lowest EEI over the last millennium is 1272–1291 ($0.4\,\mathrm{W\,m^{-2}}$), with a minimum in 1284 ($0.7\,\mathrm{W\,m^{-2}}$). This is decades after the 1257 Samalas eruption, which had an exceptional magnitude but a relatively weak climate response (Fig. S17a; Vidal et al., 2016; Guillet et al., 2017). The 20-yr period with the largest loss of OHC300 is 1803–1822 ($0.4\,\mathrm{W\,m^{-2}}$; Fig. S17d), including the 1815 Tambora eruption, and the recovery from this cluster marks the end of the LIA (Brönnimann et al., 2019). This early 1800s cluster complicates the definition of a pre-industrial baseline (e.g., Thorne et al., 2026) and the initialization of historical simulations in 1850 (Ballinger et al., 2026).
OHC anomalies in our reconstruction differ from model simulations (Figs. 10, S16, and S17). Simulated ocean cooling peaks 3 years after the eruption and recovers immediately, while the reconstructed cooling peaks at a lower amplitude. In contrast to the simulations, the reconstructed onset of recovery varies by eruption, occurring 3–10 years post-eruption and resulting in the flatter composite (Fig. 10b). The reconstructed response is likely smoothed by biological memory, or persistence, in TRW proxies, which form the bulk of our network (Fritts, 1976; Esper et al., 2015; Lücke et al., 2019; Zhu et al., 2020). This results in a muted and time-lagged volcanic response, which has long been noted in the literature (e.g., Frank et al., 2007; Anchukaitis et al., 2012; Mann et al., 2012; D’Arrigo et al., 2013; Stoffel et al., 2015; Anchukaitis and Smerdon, 2022; He et al., 2026). However, models are also known to overestimate the cooling after large-magnitude eruptions, for example, due to the way volcanic aerosols are prescribed and volcanic forcing is calculated (Timmreck et al., 2009; Timmreck, 2012; Marotzke and Forster, 2015; LeGrande et al., 2016; Marshall et al., 2025). The model–proxy discrepancy is further exacerbated by unforced variability and uncertainty in VSSI and timing (Zanchettin et al., 2019; Lücke et al., 2023), although progress has been made in resolving large eruptions in ice cores (Burke et al., 2019) and reconciling simulated and reconstructed volcanism (Stoffel et al., 2015; Zhu et al., 2020). Over 1960–2000, the timescale and magnitude of OHC loss in our reconstruction agree with instrumental datasets for three large eruptions (Fig. 4).
The loss of OHC is driven by a negative EEI at the TOA, peaking at $0.2\,\mathrm{W\,m^{-2}}$ in the composite, and arising from increased RSR (Fig. 10c,d). This is much weaker than expected for eruptions larger than Mt. Pinatubo in 1991, which had a peak EEI of around $3\,\mathrm{W\,m^{-2}}$ (Allan et al., 2014). After 3–5 years, the EEI becomes marginally positive due to reduced RSR, which is also evident in coupled model simulations, although at much larger magnitudes (Fig. S16c). By Eq. (1), the difference between dOHC300/dt $\approx C(dT/dt)$ and the EEI may be due to heat uptake $\gamma T$ by the deeper ocean. However, the difference here is likely due to the greatly underestimated volcanic forcing. The ocean cooling dOHC300/dt, with a peak of $0.4\,\mathrm{W\,m^{-2}}$ in the composite, may be a better estimate of the volcanic forcing than the reconstructed EEI.
Previous modeling studies have found expanded sea ice persisting for more than a century following large eruptions due to sea-ice–ocean feedbacks in the Arctic Ocean and North Atlantic (Zhong et al., 2011; Miller et al., 2012; Slawinska and Robock, 2018). In contrast, in our volcanic composite, the sea-ice area in both hemispheres returns to the pre-eruption value after 15–20 years (Fig. S15c,d). However, Zhong et al. (2011) find that these sea ice changes following volcanic eruptions depend on the initial condition of the coupled atmosphere–ocean–sea-ice system. Centennial sea-ice expansion that sustains cold anomalies could thus be important for some eruptions and could have played a role in initiating the LIA (Miller et al., 2012).
5.5 Comparison with last-millennium simulations
The majority of CMIP6 models do not show the millennial-scale cooling trend found in proxy reconstructions (Fig. 8 and Section 44.4; PAGES 2k Consortium, 2019), in contrast to CMIP5 models (Atwood et al., 2016). The lack of a global-mean cooling trend over 850–1850 in models is the main reason for the difference in OHC and TOA radiation with our reconstruction. If we assume that trends of all fields scale linearly with the global-mean temperature trend and adjust the model trends to match the reconstructed temperature trend, agreement improves substantially, particularly for the global means (not shown). The benefit of DA is thus that variability on seasonal to millennial timescales is directly constrained to follow the climate trajectory as sampled by proxies. Model biases may be inherited by the LIMs and thus the reconstruction, which is partly mediated by our multi-model reconstruction.
Our reconstruction provides a new dataset to compare with energy budgets simulated in coupled models. First, since the annual-mean cooling trend is likely volcanically forced (Schurer et al., 2014; Büntgen et al., 2020; Wanner et al., 2022), the lack of a simulated trend may imply that the coupled climate system response to volcanic eruptions, such as post-eruption OHC loss, is not well resolved in current models (e.g., Church et al., 2005). Other model issues that can lead to discrepancies with proxy reconstructions include uncertainty in the forcing dataset and in the aerosol representation (Lücke et al., 2023; Marshall et al., 2022; Timmreck, 2012). A false warming trend can also arise from impulsive positive radiative forcing when the prescribed background aerosols from the spin-up run are removed at the start of a transient last-millennium simulation (Gregory et al., 2013; Fyfe et al., 2021). Finally, the initialization of coupled model simulations from equilibrium is likely a poor approximation of the coupled climate system. This critique applies to both historical simulations starting in 1850, which are preceded by strong volcanism and the last-millennium cooling trend (Ballinger et al., 2026; Gebbie and Huybers, 2019), and to last-millennium simulations starting in 850, when the deep ocean is potentially still adjusting to the last deglaciation (Zhu et al., 2019). Proxy-constrained estimates of the pre-industrial deep ocean may help to provide initial conditions that better reflect earlier climate variability.
6 Conclusion
Earth’s energy imbalance is a fundamental climate metric, but our understanding of its variability on long timescales is limited by a lack of long records and large uncertainty in coupled climate models. Our reconstruction over the last millennium extends the observational record by 1000 years, providing context for energy budget variability in the absence of anthropogenic forcing. By combining proxy data with a simplified climate model using data assimilation, we obtain a consistent dataset of global surface temperatures, sea ice, top-of-atmosphere radiation, and upper-ocean heat content. In contrast to climate models, the variability on all timescales is directly constrained by proxies to follow the observed climate trajectory. The reconstruction can be used to investigate seasonal to multidecadal variability, millennial trends, and episodic events like volcanic eruptions.
We demonstrate the feasibility of reconstructing top-of-atmosphere radiation from temperature-sensitive proxies, even though the quantification of the energy imbalance and ocean heat content remains challenging during the satellite period (Hakuba et al., 2024). Through validation in pseudoproxy experiments and against satellite observations, we find that the global-mean energy imbalance, reflected shortwave radiation, and outgoing longwave radiation can be skillfully reconstructed; there is less skill in spatial patterns, especially outside of the tropics. Volcanic forcing events and a subsequent reduction in ocean heat content are captured, but with greatly underestimated magnitude. Since the simplified climate model we use is trained on last-millennium simulations, our reconstruction resolves natural forcing and the radiative response, but not anthropogenic forcing. Compared to the reconstruction of the last millennium by Gebbie and Huybers (2019), we find agreement on the magnitude of sustained energy loss, but with large error bars around zero and uncorrelated multidecadal variability, likely due to different proxy networks and reconstruction methods. These uncertainties could potentially be be reduced by including constraints on the deep ocean from proxies, either in situ (e.g., Waelbroeck et al., 2002; Lear et al., 2000) or from globally integrating noble gases (Bereiter et al., 2018a). These proxies have proven useful for reconstructions of the last deglaciation (Bereiter et al., 2018b; Baggenstos et al., 2019; Shackleton et al., 2023) and the Pleistocene (Shackleton et al., 2026).
Our reconstruction provides context for the energy gain in recent decades, which we find to be unprecedented in the context of natural, pre-industrial variability. It also reaffirms the role of volcanism in the last-millennium cooling trend, which may have been continuously forced by volcanic clusters causing an accumulation of heat loss. The discrepancy of our reconstruction with coupled simulations may point to deficiencies in the energy budget representation in the same models used for future climate projections.
Acknowledgements
We thank Jess Tierney, Sylvia Dee, and an anonymous reviewer for their comments in review, which greatly improved this manuscript. We thank Kyle Armour, Zilu Meng, Eric Mei, Aaron Donohoe, and Vince Cooper for helpful conversations. Anna Black’s master’s thesis provided preliminary results that helped motivate this study. This material is based upon work supported by the National Science Foundation under Award No. 2202526. We would like to acknowledge computing support from the Casper system (https://ncar.pub/casper) provided by the NSF National Center for Atmospheric Research (NCAR), sponsored by the National Science Foundation. This work made use of pyleoclim (Khider et al., 2022) and climlab (E. J. Rose, 2018).
Code and data availability
Code is publicly available at https://github.com/DominikStiller/energy-budget-from-proxies. The reconstruction data will be published once this manuscript has been accepted.
Copyright notice
This work has been submitted to Journal of Climate. Copyright in this work may be transferred without further notice.